# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v1

@@ Abstract
Recent studies show that Large Language Models (LLMs) achieve strong reasoning capabilities through supervised fine-tuning or reinforcement learning. However, a key approach, the Process Reward Model (PRM), suffers from reward hacking, making it unreliable in identifying the best intermediate steps. In this paper, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM performs better in assessing reasoning coherence and self-reflection, particularly when the previous reasoning step is incorrect. Furthermore, to address the inefficiency of autonomous generating PRM training data via Monte Carlo Tree Search (MCTS), we introduce a lightweight and effective data augmentation strategy called Hierarchical Node Compression (HNC) based on node merging (combining two consecutive reasoning steps into one step) in the tree structure. This approach diversifies MCTS results for HRM with negligible computational overhead, enhancing label robustness by introducing noise. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K confirm HRM’s superior generalization and robustness across diverse reasoning tasks. The code for all experiments will be released at https://github.com/tengwang0318/hierarchial_reward_model.
@@ 1 Introduction
As the scale of parameters in Large Language Models (LLMs) continues to grow [1, 2, 3, 4], their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation [5]. However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical and logical reasoning—remain an open question [6, 7, 8, 9, 10, 11, 12]. Recent approaches, such as Chain-of-Thought (CoT)[13] and Tree-of-Thought (ToT)[14], have significantly enhanced reasoning performance. Despite these advancements, CoT lacks a mechanism to halt reasoning when an intermediate step is incorrect, leading to error propagation. Meanwhile, ToT does not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
@@
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs more effectively. There are two primary approaches to enhancing the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM)[7, 8, 15, 16] and the Process Reward Model (PRM)[7, 8]. Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, while PRM is prone to reward hacking and incurs high costs for reasoning process annotation.
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). Unlike PRM, which evaluates reasoning at a fine-grained step level, HRM considers both fine-grained and coarse-grained reasoning, enabling the reward model to reflect on its own judgments and assess multi-step reasoning coherence. Traditional PRM penalizes an incorrect step without accounting for potential corrections in subsequent reasoning. In contrast, HRM enables the reward model to identify and incorporate later steps that rectify earlier errors, resulting in a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM. The PRM800K [7] dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct [4] as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results show that HRM is the most robust to reward hacking, with accuracy stabilizing at 80% as N increases. In contrast, PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as N grows.
To fully exploit the capabilities of MCTS, we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM-800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM-800K and exhibits strong generalization across other domains, including GSM8K [6] and MATH500 [7] dataset, outperforming PRM in robustness and consistency.
Our main contribution are as follows:
  • We propose the Hierarchical Reward Model (HRM), which enhances the reward model’s multi-step reasoning coherence and self-reflection by evaluating not only fine-grained reasoning steps but also coarse-grained reasoning trajectories. We validate HRM’s robustness on the PRM800K dataset using manually annotated data.
  • We introduce Hierarchical Node Compression (HNC) in MCTS for autonomous annotation, enhancing reward model training with more diverse reasoning data while incorporating controlled noise to improve score robustness. Additionally, by filtering high-quality reasoning trajectories from MCTS, we refine the policy model through fine-tuning, further enhancing its reasoning performance.
  • We investigate the generalization of HRM trained on PRM800K using auto-labeled reasoning processes from MCTS and HNC. Experimental results demonstrate that HRM achieves superior reasoning consistency and generalizes effectively across GSM8K and MATH500, outperforming PRM.
@@ 2 Related Work
Figure 1: Illustration of how ORM, PRM, and HRM handle reasoning processes. ORM evaluates the entire reasoning chain, PRM assesses individual steps but stops at errors, and HRM considers multiple consecutive steps, enabling error correction.
@@ 3.1 Hierarchical Reward Model
PRM [7, 8] focuses on fine-grained, step-wise reasoning, while ORM [7, 8] evaluates reasoning as a whole. To integrate the strengths of both, we propose the Hierarchical Reward Model (HRM), which evaluates individual reasoning steps while also ensuring multi-step coherence by analyzing consecutive steps. The training dataset for HRM consists of consecutive reasoning sequences spanning from step 1 to $N$, as illustrated in Fig. 1 and Section 4.1.
@@ 3.2 Hierarchical Node Compression in MCTS
Although process supervision enhances the reasoning capabilities of policy models, the cost of human-annotated supervision is prohibitively high [7, 8]. To address this, autonomous annotation methods based on MCTS have been proposed [10, 19]. Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct steps in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as ensuring the convergence of intermediate reasoning step scores requires a sufficiently deep and wide search tree; otherwise, the estimates remain biased. This exponential growth in complexity makes large-scale implementation challenging.
@@
As illustrated in Fig. 3, HNC assumes that each node has a sufficiently large number of child nodes. By randomly removing or merging consecutive nodes, it introduces controlled noise, enhancing the robustness of MCTS-based scoring. For instance, if each node initially has five child nodes contributing equally (20%) to the parent node’s score, merging one child node into a new branch redistributes the weights among the remaining four, increasing their contributions to 25%. This controlled perturbation improves the resilience of the scoring mechanism while expanding the dataset with minimal computational cost.
@@ 3.3 Self-Training
Without proper KL regularization or with an insufficiently weighted KL loss (i.e., a very small $\lambda$), the KL divergence grows unbounded during training. Specifically, KL loss typically ranges from 00 to $20000$, whereas the causal LM loss remains within 00 to $12$, leading to a severe loss imbalance. This causes the optimization process to excessively minimize KL divergence at the expense of task-specific reasoning performance.
To address this, we apply a logarithmic scaling to $D_{\text{KL}}(P||Q)$, stabilizing the loss landscape and ensuring a balanced trade-off between preserving general language capabilities and enhancing reasoning ability. Further details are provided in Section 4.3.
@@ 4.1 HRM
Table 2: Labeling strategy for constructing the HRM training dataset from manual annotations in PRM800K. PRM800K contains three label types: Positive, Negative, and Neutral. HRM extends PRM by incorporating multi-step reasoning.
@@
PRM 0.7 0.644 0.611 0.588 0.577
HRM 0.722 0.711 0.744 0.8 0.8
Table 3: Accuracy of Qwen2.5-72B-Math-Instruct [4] evaluated under the best-of-$N$ strategy using ORM, PRM, and HRM, which are fine-tuned on manually labeled PRM800K data.
@@
We fine-tune Qwen2.5-1.5B-Math [4] as the reward model (RM) for classifying current step as correct or incorrect. Given an input reasoning step, RM predicts logits for the positive and negative classes, denoted as $l_{\text{positive}}$ and $l_{\text{negative}}$, respectively. The confidence score is obtained by applying the softmax function:
@@
where $P(y=\text{positive}\mid x)$ denotes the probability of current step being correct. This probability serves as the reward score assigned by RM. Detailed training information is provided in Appendix A.1.
To evaluate the performance of ORM, PRM, and HRM, we employ Qwen2.5-72B-Math-Instruct [4] as the policy model and implement the best-of-$N$ strategy. Specifically, ORM selects the best result from $N$ complete reasoning trajectories, while PRM and HRM score $N$ intermediate reasoning steps and select the most promising one at each step. For PRM and HRM, we consider the completion of a formula as an intermediate reasoning step, enabling a finer-grained evaluation mechanism. Table 3 presents the results, showing that ORM and PRM exhibit significant fluctuations, with accuracy decreasing as $N$ increases. In contrast, HRM maintains stable performance, converging to an accuracy of 80% as $N$ grows. Detailed information is provided in Appendix A.1
@@ 4.2 HNC
Table 4: Accuracy of PRM and HRM under different policy models using the best-of-$N$ strategy on the PRM800K dataset. The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct.
Table 4 presents the accuracy results of various policy models under PRM and HRM settings on the PRM800K dataset. Although both PRM and HRM training data are derived from MCTS with Qwen2.5-7B-Math, we evaluate the fine-tuned models using different policy models, where HRM consistently demonstrates greater stability and robustness compared to PRM.
The relatively lower performance of Qwen2.5-72B-Math-Instruct can be attributed to the tree height constraints imposed by MCTS, which necessitate answer generation within a predefined template and a fixed number of steps. Although Qwen2.5-72B-Math-Instruct demonstrates strong reasoning capabilities, its highly structured training process renders it susceptible to performance degradation when deviating from its learned format. To ensure computational feasibility, our MCTS framework enforces limitations on tree depth, which further amplifies this effect.
@@ 4.3 Self-Training
We adapt the method described in Section 3.3 to filter high-quality reasoning data and train the policy model. Fig. 4 illustrates that when $\lambda$ is small (e.g., 0.001), the fine-tuned model rapidly loses its generalization ability within just a few iterations, causing the KL loss to escalate to approximately 20,000. In contrast, the causal LM loss remains within the range of 0 to 12, leading to a significant imbalance. This discrepancy underscores the necessity of applying logarithmic scaling to the KL term in the objective function, as discussed in Section 3.3. Conversely, when $\lambda$ is excessively large (e.g., 10.0), the model prioritizes adherence to the reference distribution, resulting in slower convergence and constrained improvements in reasoning capability.
We perform SFT using high-MC-score reasoning data from the PRM/HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on high-MC-score reasoning data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 5 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.

### v2

@@ Abstract
Recent studies show that Large Language Models (LLMs) achieve strong reasoning capabilities through supervised fine-tuning or reinforcement learning. However, a key approach, the Process Reward Model (PRM), suffers from reward hacking, making it unreliable in identifying the best intermediate step. In this paper, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM excels in evaluating multi-step reasoning coherence and self-reflection, especially in scenarios where the previous reasoning step is incorrect but the subsequent step successfully identifies and corrects the error. Furthermore, to address the inefficiency of autonomous annotating PRM training data via Monte Carlo Tree Search (MCTS), we introduce a lightweight and effective data augmentation strategy called Hierarchical Node Compression (HNC) based on node merging (combining two consecutive reasoning steps into one step) in the tree structure. By applying HNC to MCTS-generated reasoning trajectories, we enhance the diversity and robustness of HRM training data while introducing controlled noise with minimal computational overhead. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K dataset confirm HRM’s superior generalization and robustness across diverse reasoning tasks.
@@ 1 Introduction
As the scale of parameters in Large Language Models (LLMs) continues to grow [1, 2, 3, 4], their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation [5]. However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical and logical reasoning—remain an open question [6, 7, 8, 9, 10, 11, 12, 13]. Recent approaches, such as Chain-of-Thought (CoT)[14] and Tree-of-Thought (ToT)[15], have significantly enhanced reasoning performance. Despite these advancements, CoT lacks a mechanism to halt reasoning when an intermediate step is incorrect, leading to error propagation. Meanwhile, ToT does not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
@@
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs more effectively. There are two primary approaches to enhancing the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM) [7, 8, 16, 17] and the Process Reward Model (PRM) [7, 8]. Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, making it difficult to pinpoint which reasoning steps contribute to the final answer [7, 8]. PRM, in contrast, provides finer-grained supervision by evaluating reasoning step by step. However, it is prone to reward hacking [18, 19, 20], where models exploit reward signals rather than genuinely improving reasoning, undermining reliability in complex tasks. Moreover, the high annotation cost associated with PRM makes large-scale deployment challenging.
@@
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). Traditional PRM penalizes an incorrect step without considering potential corrections in subsequent reasoning. In contrast, HRM assesses reasoning coherence across multiple steps, allowing the reward model to identify and incorporate later steps that rectify earlier errors, leading to a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM. The PRM800K [7] dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct [21] as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results (as shown in Table 3) demonstrate that HRM is the most robust to reward hacking. The policy model with HRM maintains stable performance, with accuracy stabilizing at 80% as $N$ increases. In contrast, policy models with PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as $N$ grows.
To fully exploit the capabilities of MCTS, we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM800K dataset and exhibits strong generalization across other domains, including GSM8K [6] and MATH500 [7] dataset, outperforming PRM in robustness and consistency.
Our main contribution are as follows:
  • We propose the Hierarchical Reward Model (HRM), which enhances multi-step reasoning coherence and self-reflection by evaluating both fine-grained reasoning steps and coarse-grained reasoning trajectories. We validate HRM’s robustness on the PRM800K dataset using manually annotated data.
  • We introduce Hierarchical Node Compression (HNC) in MCTS for autonomous annotation, refining existing reasoning trajectories by generating additional step-wise labels with controlled noise to improve score robustness. We train HRM on PRM800K dataset with auto-annotated reasoning data from HNC and MCTS, demonstrating its improved robustness over PRM. Furthermore, HRM trained on PRM800K dataset exhibits superior reasoning consistency and generalization across GSM8K and MATH500 dataset, outperforming PRM.
  • Additionally, we enhance the policy model through fine-tuning on high-quality reasoning trajectories filtered from MCTS, further improving its reasoning performance and robustness.
@@ 2.3 PRM
Figure 1: Illustration of how ORM, PRM, and HRM handle reasoning processes. ORM evaluates the entire reasoning chain, PRM assesses individual steps but stops at errors, and HRM considers multiple consecutive steps, enabling error correction. The figure also demonstrates how HRM constructs its training dataset by merging two consecutive steps.
Figure 2: Illustration of the MCTS-based automated reasoning annotation process. The left side depicts a tree structure where each node represents a reasoning step, simulated using the Tree-of-Thought approach with MCTS. The right side visualizes the assigned scores for each step in the reasoning tree.
@@ 3.1 Hierarchical Reward Model
PRM [7, 8] focuses on fine-grained, step-wise reasoning, while ORM [7, 8] evaluates reasoning as a whole. To integrate the strengths of both, we propose the Hierarchical Reward Model (HRM), which evaluates individual reasoning steps while also ensuring multi-step coherence by analyzing consecutive steps. The training dataset for HRM consists of consecutive reasoning sequences spanning from step 1 to $N$, as illustrated in Fig. 1 and Section 4.1. HRM training data is a superset of PRM training data, constructed by merging consecutive reasoning steps. Specifically, while PRM training data only contains individual steps with their respective reward scores, HRM training data additionally incorporates merged step pairs to enhance reasoning coherence and improve self-reflection in error correction. Formally, let $D$ represent the training dataset, $N$ denote the total number of reasoning steps in a sequence, $s_{i}$ be the $i$-th reasoning step, and $R(\cdot)$ be the reward function that assigns a score to a step. The training datasets for PRM and HRM are defined as:
$D_{\text{PRM}}=\{(s_{i},R(s_{i}))\mid 1\leq i\leq N\},$
$D_{\text{HRM}}=D_{\text{PRM}}\cup\{(s_{i}+s_{i+1},R(s_{i}+s_{i+1}))\mid 1\leq i<N\}.$
@@
Although HRM training data incorporates merged reasoning steps, the model remains step-wise in inference, assigning a reward based solely on the current step $s_{i}$, similar to PRM.
@@ 3.2 Hierarchical Node Compression in MCTS
Although process supervision enhances the reasoning capabilities of policy models, the cost of human-annotated supervision is prohibitively high [7, 8]. To address this, autonomous annotation methods based on MCTS have been proposed [10, 24]. Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct steps in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as achieving reliable estimates of intermediate reasoning step scores requires a sufficiently deep and wide search tree to reduce variance and mitigate bias; otherwise, the estimates may remain unreliable. This exponential growth in complexity makes large-scale implementation challenging.
@@
As illustrated in Fig. 3, HNC assumes that each node has a sufficiently large number of child nodes. By randomly removing or merging consecutive nodes, it introduces controlled noise, enhancing the robustness of MCTS-based scoring. Before HNC, each child node contributes $\frac{1}{N}$ to the total score. After HNC removes a random node, the remaining child nodes redistribute their weights to $\frac{1}{N-1}$, increasing their individual influence. Since child nodes are independent and identically distributed (i.i.d.) from the parent’s perspective, the expectation of the parent score remains unchanged. However, the variance increases from $\frac{\sigma^{2}}{N}$ to $\frac{\sigma^{2}}{N-1}$, introducing controlled noise that enables data augmentation at an extremely low computational cost. When $N$ is sufficiently large, this variance change remains moderate while still facilitating effective data augmentation.
@@ 3.3 Self-Training
Without proper KL regularization or with an insufficiently weighted KL loss (i.e., a very small $\lambda$), the KL divergence grows unbounded during training. Specifically, KL loss typically ranges from 00 to $20000$, whereas the causal LM loss remains within 00 to $12$, leading to a severe loss imbalance. This causes the optimization process to excessively minimize KL divergence at the expense of task-specific reasoning performance. To address this, we apply a logarithmic scaling to $D_{\text{KL}}(P||Q)$, stabilizing the loss landscape and ensuring a balanced trade-off between preserving general language capabilities and enhancing reasoning ability. Further details are provided in Section 4.3.
@@ 4.1 HRM
Table 2: Labeling strategy for constructing the HRM training dataset from manual annotations in PRM800K dataset. PRM800K dataset contains three label types: Positive, Negative, and Neutral. HRM extends PRM by incorporating multi-step reasoning.
@@
PRM 0.700 0.644 0.611 0.588 0.577
HRM 0.722 0.711 0.744 0.800 0.800
Table 3: Accuracy of Qwen2.5-72B-Math-Instruct [21] evaluated under the best-of-$N$ strategy using ORM, PRM, and HRM on the PRM800K test set. ORM, PRM, and HRM are fine-tuned on the manually labeled PRM800K dataset.
@@
We fine-tune Qwen2.5-1.5B-Math [21] as the reward model (RM) for classifying given reasoning step as correct or incorrect. Given an input, RM predicts logits for the positive and negative classes, denoted as $l_{\text{positive}}$ and $l_{\text{negative}}$, respectively. The confidence score is obtained by applying the softmax function:
@@
where $P(y=\text{positive}\mid x)$ denotes the probability of given reasoning step being correct. This probability serves as the reward assigned by RM. Detailed information is provided in Appendix A.1.
To evaluate the performance of ORM, PRM, and HRM, we employ Qwen2.5-72B-Math-Instruct [21] as the policy model and implement the best-of-$N$ strategy. Specifically, ORM selects the best result from $N$ complete reasoning trajectories, while PRM and HRM score $N$ intermediate reasoning steps and select the most promising one at each step. For PRM and HRM, we consider the completion of a formula as an intermediate reasoning step, enabling a finer-grained evaluation mechanism. Table 3 presents the results, showing that the accuracy of the policy model with ORM and PRM exhibits significant fluctuations, decreasing as $N$ increases. In contrast, the policy model with HRM maintains stable performance, converging to an accuracy of 80% as $N$ grows.
@@ 4.2 HNC
Table 4: Accuracy of different policy models under PRM and HRM using the best-of-$N$ strategy on the PRM800K test set [7]. The training data for both PRM and HRM are derived from MCTS with Qwen2.5-7B-Math-Instruct [21]
Table 4 presents the accuracy results of various policy models under PRM and HRM settings on the PRM800K dataset. Although both PRM and HRM training data are derived from MCTS with Qwen2.5-7B-Math-Instruct, we evaluate HRM and PRM using different policy models, where HRM consistently demonstrates greater stability and robustness compared to PRM.
The relatively lower performance of Qwen2.5-72B-Math-Instruct (as shown in Table 4) can be attributed to the tree height constraints imposed by MCTS, which require answer generation to follow a predefined template within a fixed number of steps. While Qwen2.5-72B-Math-Instruct exhibits strong reasoning capabilities [21], its highly structured training process makes it more vulnerable to performance degradation when deviating from its learned format.
@@ 4.3 Self-Training
We adapt the method described in Section 3.3 to filter high-quality reasoning data and train the policy model. Fig. 4 illustrates that when $\lambda$ is small (e.g., 0.001), the fine-tuned policy model rapidly loses its generalization ability within just a few iterations, causing the KL loss to escalate to approximately 20,000. In contrast, the causal LM loss remains within the range of 0 to 12, leading to a significant imbalance. This discrepancy underscores the necessity of applying logarithmic scaling to the KL term in the objective function, as discussed in Section 3.3. Conversely, when $\lambda$ is excessively large (e.g., 10.0), the model prioritizes adherence to the reference distribution, resulting in slower convergence and constrained improvements in reasoning capability.
We fine-tune Qwen2.5-7B-Math-Instruct [21] using high-MC-score reasoning data extracted from the PRM and HRM training datasets. Qwen-7B-HRM denotes the policy model fine-tuned on high-MC-score reasoning data from HRM’s training set, while Qwen-7B-PRM follows the same procedure for PRM’s training set. We set $\lambda$ to 0.5. Table 5 further validates that SFT enhances the policy model’s reasoning capability by leveraging high-quality data, with HRM demonstrating greater robustness compared to PRM.